# CollabStory: Multi-LLM Collaborative Story Generation and Authorship Analysis

### v2

@@ Abstract
The rise of unifying frameworks that enable seamless interoperability of Large Language Models (LLMs) has made LLM-LLM collaboration for open-ended tasks a possibility. Despite this, there have not been efforts to explore such collaborative writing. We take the next step beyond human-LLM collaboration to explore this multi-LLM scenario by generating the first exclusively LLM-generated collaborative stories dataset called CollabStory. We focus on single-author ($N=1$) to multi-author (up to $N=5$) scenarios, where multiple LLMs co-author stories. We generate over 32k stories using open-source instruction-tuned LLMs. Further, we take inspiration from the PAN tasks Bevendorff et al. (2023) that have set the standard for human-human multi-author writing tasks and analysis. We extend their authorship-related tasks for multi-LLM settings and present baselines for LLM-LLM collaboration. We find that current baselines are not able to handle this emerging scenario. Thus, CollabStory is a resource that could help propel an understanding as well as the development of techniques to discern the use of multiple LLMs. This is crucial to study in the context of writing tasks since LLM-LLM collaboration could potentially overwhelm ongoing challenges related to plagiarism detection, credit assignment, maintaining academic integrity in educational settings, and addressing copyright infringement concerns. We make our dataset and code available at https://anonymous.4open.science/r/multi_llm_story_writing-55AE/README.md.
@@ 1 Introduction
Collaborative creative story writing entails multiple authors contributing separate segments to form a coherent storyline (see Figure 1 for our dataset schema). Although individual LLMs excel at generating story plots, collaborative writing presents unique hurdles. Models must seamlessly continue the existing storylines generated so far by other models, even if they do not align perfectly with their own language distribution. The rise of multi-agent Artificial Intelligence (AI) underscores the potential for combining the expertise of agents specialized in various tasks. While previous mixture-of-experts scenarios focused on agents proficient in task-oriented settings Zhang et al. (2024); Pan et al. (2024); Liu et al. (2023); Li et al. (2023b, a), the emergence of LLMs conversing for continuous generative tasks in open domains is noteworthy. Imagine the possibilities when multiple LLMs collaborate; one LLM can generate compelling stories, but what if we put them together?
@@
We focus on single-author ($N=1$) to multi-author (up to $N=5$) scenarios, where multiple LLMs co-author creative stories. This exploration is novel, as previous studies have primarily focused on human-LLM collaboration. Towards this goal, we generate the first multi-LLM collaborative story dataset called CollabStory using open-source LLMs. We select 5 frequently used LLMs (with number of downloads on HuggingFace for May 2024 provided in parenthesis): Meta’s Llama ($>540k$ downloads, Touvron et al. (2023)), Mistral.ai’s Mistral ($>1000k$ downloads, Jiang et al. (2023)), Google’s Gemma ($>180k$ downloads, Team et al. (2024)), AllenAI’s Olmo ($>26k$ downloads, Groeneveld et al. (2024)) and Microsoft’s Orca ($>22k$ downloads, Mukherjee et al. (2023)) to replicate a scenario in which commonly used LLMs from different organizations are being used in conjunction towards a single task. We demonstrate how one such dataset can be developed and the considerations involved in building an iterative Multi-LLM story-writer. We present an evaluation of the generated texts and their quality as the number of authors increases from one to five. Since this is the first such multi-LLM dataset, our evaluation provides insights into the current state of LLMs and their ability to collaborate on open-ended generation tasks. We then take inspiration from the PAN tasks Bevendorff et al. (2023) that have set the standard for multi-author writing tasks and analysis for human-human collaboration for over 15 years. We adapt their authorship-related task suite to the machine-machine settings for problems such as authorship verification and attribution, and demonstrate that current baselines are challenged by this emerging scenario. CollabStory is the first resource that could help propel an understanding as well as the development of new techniques to discern the use of multiple LLMs in text.
Our work is motivated by the implications of Multi-LLM settings for different stakeholders (LLM developers, end-users) and considerations (such as credit assignment, legality of usage) arising in the generative AI landscape. As an example, a malicious actor might assemble texts from different LLMs together in one document to evade attribution and successfully spread misinformation. Our findings elaborate on the tasks our dataset enables and emphasize the importance of tackling the incoming challenges of machine-machine collaboration.
@@ 3.2 LLM prompting
Mistral Gemma Llama Olmo Orca
Years ago, in the depths of a dark and desolate forest, I made a promise to a witch. Desperate for her help to save my beloved firstborn son, I agreed to offer her my firstborn in return for her healing magic. The old crone’s cackling laughter still echoed in my ears as I made that fateful vow. But time passed, and my wife bore us three more children. With each birth, I was filled with relief and gratitude, knowing that our family had been spared the tragedy that had once threatened to tear us apart. Yet no matter how hard I tried, I couldn’t bring myself to fulfill my promise to the witch. Now, years have passed, and the witch has grown tired of waiting. Her messages, delivered by raven messenger, are becoming more and more urgent. The ground beneath my feet seems to shake with her anger, and I can’t help but fear for the consequences of my broken vow. Desperate for a solution, I turn to the only person who might be able to help me. But will she be able to save me from the wrath of the vengeful witch? Her pale, ethereal face held a cruel smile, her voice like the rustling of dry leaves in a desolate forest. "Your son’s life hangs in the balance, dear one," she hissed, her voice dripping with malice. "Yet, you have failed to fulfill your part of the bargain." Panic surged through me as I realized the witch was right. I had never presented her with a child. The burden of my desperation had become a heavier chains, tethering me to an unspoken debt. Desperation fueled the flames of my fear, as I knew that if I failed, not only would my son succumb to his fate, but the wrath of the vengeful witch would descend upon me. Time seemed to warp, stretching into an agonizing eternity as I considered my options. Could I find a way to appease her, to escape the consequences of my foolish decision? I paced back and forth in my chambers, my mind racing with thoughts of the witch and the debt I owed her. I had been so desperate to save my son, I had agreed to give her my firstborn child in exchange for her healing magic. Now, years had passed and the witch had grown tired of waiting. She had come to me, demanding that I fulfill my end of the bargain. I knew that I could not give her my son. He was all I had left of my beloved husband, and I would rather die than lose him. But I also knew that I could not defy the witch. She had the power to take my life, and the lives of my children, with just a wave of her hand. I spent hours pacing, trying to think of a way to appease her. But no matter what I did, I could not come up with a solution. I was trapped, and I knew it. The witch had me in her grasp, and she would not let me go. Finally, in desperation, I decided to seek out the help of a wise old woman who lived in the village. A few days ago, I came to her, my heart heavy with grief and my soul wounded. She, in turn, looked at me with a kind and gentle expression. Then, with a wave of her hand, she began her spell. I watched as the darkness in my heart began to fade, and my love for my children grew stronger than ever. She whispered words of power and love, and the bond between my wife and I was restored. From that day forward, we were a family once more, united by the magic of the wise old woman. The wise old woman had always been a renowned healer in our village. Her knowledge of herbs and potions was unmatched, and many came to her seeking relief from various ailments. I was one of those who sought her help, when my firstborn son fell gravely ill. His fever raged, and his tiny body was wracked with pain. Desperate for her help to save my beloved firstborn son, I agreed to give her my firstborn in exchange for her healing magic. I returned to my wife, heartbroken and ashamed, to tell her of the bargain I had made. But time passed, and my wife bore us three more children. With each birth, I was filled with relief and gratitude, knowing that our family had been spared the tragedy of losing a child. The wise old woman had honored her part of the bargain, and my son grew strong and healthy under her care. He became a skilled healer in his own right, following in the footsteps of the wise old woman who had saved his life.
Table 4: Example of a 5-part LLM story from CollabStory dataset for the prompt: “Years ago, you promised your firstborn to a witch. Since then, despite your best efforts, you can’t seem to give him away. The witch is starting to get pretty mad.
@@ 4 Dataset Analysis
We compare the LLM-generated single and multi-authored texts with the single-authored human written stories to study the relative quality of LLM generated stories as the number of authors present increases. To do this, we report the average and standard deviation of different measures such as the average number of (i) words, (ii) sentences, (iii) vocabulary richness, (iv) percentage of stopwords, (v) readability scores, (vi) entropy, and (vii) coherence scores using the TextDescriptives Library [^7] https://github.com/HLasse/TextDescriptives for different numbers of authors ($N$) in Figure 2. We measure (vii) creativity using OCSAI (Open Creativity Scoring with Artificial Intelligence) Organisciak et al. (2023) that provides text creativity scores using an LLM (GPT-4) fine-tuned on human annotations of creativity. From Figure 2, we see that across most measures, there isn’t a significant deviation or decline as the number of authors is increased, i.e. collaboration or introducing multiple LLMs in the generation setting does not disturb the quality of the stories generated. In fact, the only statistically significant difference between LLM-coauthored and human-written stories in vocabulary richness measured by type-token-ratio (TTR) for $N\in[2,3,4]$. For all other measures, our dataset’s machine-generated stories follow similar distributions to human-written stories. Automated measures of readability and coherence utilized the human-generated stories as the reference text, and even for these measures, we see that collaboratively generated text scores do not deviate from the reference score distribution. This analysis indicates that LLMs are able to collaboratively generate stories without any significant changes in quality through sequential prompting. Detailed statistics for each LLM are provided in the Appendix (Table 7).
Figure 4: To evaluate the continuity of story parts, we sample two potential continuations to a story part, one correct and one incorrect. In this figure, we see an example of a story part on the left and two candidates to be evaluated for continuity. In green are some terms highlighted to show continuity in topics, characters, and plot lines, and in red are highlighted sections of the text that suggest discontinuity.
@@ 4.1 Story Continuity
An important aspect of collaborative storywriting is the notion of continuity. Particularly, we wanted to measure if the different story parts generated sequentially by different LLM authors followed a logical and cohesive plot. To evaluate this, we used a prompt-based evaluation using GPT-4o OpenAI (2024) to discern if two consecutive story parts continue a story well or not. To do this, we develop two types of pairs of story parts: one pair containing the "Correct" or the positive sample which points to the actual next part in the story, or "Incorrect" which is the negative sample that is either a random story part drawn from a different story or from the same one (but not the true next one, any other part). In other words, for the negative samples taken from within the same story, we ensure to not use the correct consecutive part as a negative sample. An example of this setup is illustrated in Figure 4, where given a story part, there is an actual continuation ("Correct") and a random negative sample ("Incorrect") story part drawn from either the same story or a different one. We then use the following prompt structure to obtain continuity evaluations for both candidate next parts and then compare if the actual continuation was the one with a better continuation evaluation. We provide GPT-4o a pair of texts at a time (out of a total of 600 pairs) and ask it the following question: “Does Part 2 serve as a good continuation of Part 1 in terms of logical flow, coherence, and consistency? Please respond with Yes or No.” GPT-4o generated the binary "yes/no" responses and an explanation. We then compare the percentage of story parts for which either the correct answer got a "Yes" and the incorrect one got a "No" response (Correct Wins), if both candidate continuations were evaluated the same (Neither Wins), or if the Incorrect answer got a "Yes" and the Correct one got a "No". We present the results in Table 5. We see that when the negative story parts are sampled from different stories than the one that the first part belongs to, then there are 0 cases for which the incorrect part wins. This is an expected result, as story parts that belong to different stories would have different topics, plot lines, characters, and so on. Hence, detecting this discontinuity should not be hard, as is the case. An example of a case where Neither Wins and both candidate story parts get an evaluation as being discontinuous is provided in Figure 4. As illustrated by this example, it is crucial to note that our prompt encourages the model to evaluate the logical or common-sense flow of the stories. Hence, topic overlap alone was insufficient to mark two parts as continuous. In the harder setting, when the negative sample is also a story part from within the same story, we see that there is a marked increase in both the Neither Wins and Incorrect Wins scenarios (see Table 5) where Incorrect Wins goes from 0 to 9.67% and Neither Wins has an over 30% increase. This is expected since many of the story parts were not the next immediate part but a few sequences apart. They all follow the same story and share some logical flow despite the distance in their occurrences in the story. For example, the fourth part of story might get a positive continuity evaluation, just as the second, third, and fifth would if compared with the first part of the story. Over 85% of Neither Wins story pairs had a continuous evaluation for both candidates (142 out of 166). Thus, out-of-the-box LLMs are able to follow plot lines and logic even when continuing each others’ partially written parts at any point in the story.

### v3

@@ Abstract
The rise of unifying frameworks that enable seamless interoperability of Large Language Models (LLMs) has made LLM-LLM collaboration for open-ended tasks a possibility. Despite this, there have not been efforts to explore such collaborative writing. We take the next step beyond human-LLM collaboration to explore this multi-LLM scenario by generating the first exclusively LLM-generated collaborative stories dataset called CollabStory. We focus on single-author ($N=1$) to multi-author (up to $N=5$) scenarios, where multiple LLMs co-author stories. We generate over 32k stories using open-source instruction-tuned LLMs. Further, we take inspiration from the PAN tasks Bevendorff et al. (2023) that have set the standard for human-human multi-author writing tasks and analysis. We extend their authorship-related tasks for multi-LLM settings and present baselines for LLM-LLM collaboration. We find that current baselines are not able to handle this emerging scenario. Thus, CollabStory is a resource that could help propel an understanding as well as the development of techniques to discern the use of multiple LLMs. This is crucial to study in the context of writing tasks since LLM-LLM collaboration could potentially overwhelm ongoing challenges related to plagiarism detection, credit assignment, maintaining academic integrity in educational settings, and addressing copyright infringement concerns. We make our dataset and code available at https://github.com/saranya-venkatraman/CollabStory.
@@ 1 Introduction
Collaborative creative story writing entails multiple authors contributing separate segments to form a coherent storyline (see Figure 1 for our dataset schema). Although individual LLMs excel at generating story plots, collaborative writing presents unique hurdles. Models must seamlessly continue the existing storylines generated so far by other models, even if they do not align perfectly with their own language distribution. The rise of multi-agent Artificial Intelligence (AI) underscores the potential for combining the expertise of agents specialized in various tasks. While previous mixture-of-experts scenarios focused on agents proficient in task-oriented settings Zhang et al. (2024b); Pan et al. (2024); Liu et al. (2023); Li et al. (2023b, a), the emergence of LLMs conversing for continuous generative tasks in open domains is noteworthy. Imagine the possibilities when multiple LLMs collaborate; one LLM can generate compelling stories, but what if we put them together?
@@
We focus on single-author ($N=1$) to multi-author (up to $N=5$) scenarios, where multiple LLMs co-author creative stories. This exploration is novel, as previous studies have primarily focused on human-LLM collaboration. Towards this goal, we generate the first multi-LLM collaborative story dataset called CollabStory using open-source LLMs. We select 5 frequently used LLMs (with number of downloads on HuggingFace for May 2024 provided in parenthesis): Meta’s Llama ($>540k$ downloads, Touvron et al. (2023)), Mistral.ai’s Mistral ($>1000k$ downloads, Jiang et al. (2023)), Google’s Gemma ($>180k$ downloads, Team et al. (2024)), AllenAI’s Olmo ($>26k$ downloads, Groeneveld et al. (2024)) and Microsoft’s Orca ($>22k$ downloads, Mukherjee et al. (2023)) to replicate a scenario in which commonly used LLMs from different organizations are being used in conjunction towards a single task.
We demonstrate how one such dataset can be developed and the considerations involved in building an iterative Multi-LLM story-writer. We present an evaluation of the generated texts and their quality as the number of authors increases from one to five. Since this is the first such multi-LLM dataset, our evaluation provides insights into the current state of LLMs and their ability to collaborate on open-ended generation tasks. We then take inspiration from the PAN tasks Bevendorff et al. (2023) that have set the standard for multi-author writing tasks and analysis for human-human collaboration for over 15 years. We adapt their authorship-related task suite to the machine-machine settings for problems such as authorship verification and attribution, and demonstrate that current baselines are challenged by this emerging scenario. CollabStory is the first resource that could help propel an understanding as well as the development of new techniques to discern the use of multiple LLMs in text.
Our work is motivated by the implications of Multi-LLM settings for different stakeholders (LLM developers, end-users) and considerations (such as credit assignment, and legality of usage) arising in the generative AI landscape. As an example, a malicious actor might assemble texts from different LLMs together in one document to evade attribution and successfully spread misinformation. Our findings elaborate on the tasks our dataset enables and emphasize the importance of tackling the incoming challenges of machine-machine collaboration.
@@ 2 Related Work
It is interesting to consider what it means for an entity to be considered an author. LLMs do not have an innate authorship style in the way that most human writers might. It is true that an LLM likely has an emergent style that is some combination or influenced by all the different authors present in its training set. However, for the scope of this work, we turn to current literature that broadly treats authorship as a text source, hence, all the text originating from the same LLM is being treated as one group or class Soto et al. (2024); Zhang et al. (2024a); Richburg et al. (2024); Guo et al. (2024); Koike et al. (2024). For example, even in the case of human writers, the same author might shift their style based on the situational context and switch between internal personas (say, a formal tone to a comical tone). Even with such style shifts, all the text that originates from the same person would still be considered to be authored by the same entity. In a similar manner, in the absence of a prompt that forces the LLM to adhere to some strict style (such as “write a story as an 18th Century poet” or “write a story as an angry politician”), the text being generated by LLMs seem to still have some emergent and consistent patterns that enable its detection and attribution Kumarage and Liu (2023); Antoun et al. (2024); Silva and Frommholz (2023); Tripto et al. (2024). Thus, in this work, when the LLM generates text from its learned distribution in the absence of any explicit prompting that might alter style in specific ways, the text being generated is considered as being a part of the LLM’s authorship.
@@ 3.2 LLM prompting
Llama Olmo Orca Gemma Mistral
Time and time again, you failed. Today, you succeeded. The golden trophy gleamed in your hands, casting a warm glow over the crowded arena. The cheers of the crowd washed over you, a cacophony of sound that threatened to drown out your own thoughts. You couldn’t believe it. You had finally done it. You had won. But the victory was tinged with a sense of sadness. You had sacrificed so much to get here. The countless hours of training, the endless miles of running, the early morning wake-up calls and the sleepless nights. It had all been worth it, you told yourself, but a part of you couldn’t help but wonder if it was all worth it . As you held the trophy aloft, the crowd cheered even louder. You closed your eyes and let the moment wash over you. This was what it was all about. This was why you had worked so hard. This was why you had pushed yourself to the limit. This was why you had failed so many times before. But deep down, you knew that this victory was not just about you. There was something more to the silver-haired coach , something that you had seen the moment he had entered the arena. A quiet determination that had inspired you to push yourself beyond your limits. You had never believed in destiny or fate, but looking at the coach, you couldn’t help but think that maybe there was something more to the connection between the two of you. You could feel the eyes of the crowd on you, the weight of their expectations. But you were not one to shy away from a challenge. You turned to face the coach, a determined smile on your face. "Thank you for believing in me," you said, your voice quiet but steady. The coach’s face softened, and you could see the relief and pride on his face. "I never doubted your abilities, Zephyrus. You have always had the potential to be great." The words felt like a weight had been lifted off your shoulders. You knew that the road ahead would not be easy, but with the support of the coach, you felt more confident than ever. You took a step closer to the coach, reaching out to touch his hand. "I’ll do everything I can to live up to your expectations," you said, your voice trembling with emotion. The crowd roared in response , their cheers drowning out any lingering doubts you may have had. You looked out at their faces, a sea of adoration and support, and felt a surge of determination wash over you. You would not let them down, not now, when you were so close to achieving your ultimate goal. As you took your place on the podium, you couldn’t help but think back to the countless hours you had spent training, the blood, sweat, and tears you had poured into this journey. And yet, here you were, a champion, a beacon of hope for all those who had believed in you. You glanced over at your family, your friends, and your coach, their faces beaming with pride. You knew that they had sacrificed just as much as you, if not more, and you vowed to honor their faith in you by pushing yourself to new heights. With a deep breath, you stepped up to the microphone, ready to deliver your acceptance speech. You spoke from the heart, sharing your story, your struggles, and your dreams. The words tumbled out effortlessly, revealing a vulnerability that resonated with the audience. Some wept, others roared in applause, but you felt a sense of peace that transcended all the accolades and recognition. The golden trophy felt like a symbol of your journey, not your destination. It was the culmination of countless sleepless nights, tireless training sessions, and the unwavering support of those who believed in you. But the real prize, you realized, was not the gold, but the love and admiration of the people who had watched you grow into the champion you had become. You lifted the trophy high, your heart swelling with gratitude and pride. This moment, this feeling, was what made it all worth it. As the crowd roared their approval, you took a moment to look around, taking in the faces of your family, friends, and mentors. Their eyes were filled with joy and pride, and it was in that instant that you knew you had not only won a competition, but also their hearts. With a triumphant grin, you raised the trophy once more, basking in the glory of your victory, but more importantly, the love and admiration that came with it. And in that moment, you knew that no matter what challenges came your way, you would face them with the same determination and perseverance that had brought you to this point.
Table 4: Example of a 5-part LLM story from CollabStory dataset for the prompt: “Time and time again, you failed. Today, you succeeded.” As the story progresses, new characters are introduced (such as the coach highlighted in Part 2), with future parts linking to concepts from previous parts. In this example, the first part asks a question about the worth of the trophy, and the last part offers a completion to that sub-plot (highlighted in Part 1 and Part 5).
@@ 4 Dataset Analysis
We compare the LLM-generated single and multi-authored texts with the single-authored human written stories to study the relative quality of LLM generated stories as the number of authors present increases. To do this, we report the average and standard deviation of different measures such as the average number of (i) words, (ii) sentences, (iii) vocabulary richness, (iv) percentage of stopwords, (v) readability scores, (vi) entropy, and (vii) coherence scores using the TextDescriptives Library [^7] https://github.com/HLasse/TextDescriptives for different numbers of authors ($N$) in Figure 2. We measure (vii) creativity using OCSAI (Open Creativity Scoring with Artificial Intelligence) Organisciak et al. (2023) that provides text creativity scores using an LLM (GPT-4) fine-tuned on human annotations of creativity. Its high correlation with human judgments makes it a more reliable measure of overall creativity and divergent thinking than other methods that rely on semantic similarity. We used the Wilcoxon signed-rank test, a non-parametric method, to evaluate differences in text properties, including vocabulary richness, between human-written and LLM-coauthored stories (ranging from 1 to 5 LLM contributors). This test, with a significance level of p=0.05, accounts for non-normal distributions in the data.
From Figure 2, we see that across most measures, there isn’t a significant deviation or decline as the number of authors is increased, i.e. collaboration or introducing multiple LLMs in the generation setting does not disturb the quality of the stories generated. In fact, the only statistically significant difference between LLM-coauthored and human-written stories in vocabulary richness measured by type-token-ratio (TTR) for $N\in[2,3,4]$. The lower scores in vocabulary richness, coherence, or creativity should not be considered inferior performance. Higher coherence scores in LLM-generated stories often originate from the tendency of LLMs to emphasize explicit logical flow and sentence-level similarity, producing semantically aligned and structured outputs. In contrast, human stories frequently employ nuanced storytelling techniques, including implicit connections, creative leaps, and varied narrative structures Bruner (2003), which, while enriching the narrative, may lead to lower scores in metrics like embedding-based coherence.
Human authors may exhibit lower vocabulary richness because, according to the theory of linguistic individuality Nini (2023), individuals develop a mental “key profile” or preferred set of linguistic units, including specific words and phrases, that reflect their unique style and cognitive preferences. When writing stories, humans often consciously or unconsciously draw upon this limited set of vocabulary to maintain coherence, consistency, and fluency, which can restrict lexical variety compared to AI systems that leverage a broader and less constrained lexicon. For all other measures, our dataset’s machine-generated stories follow similar distributions to human-written stories. Automated measures of readability and coherence utilized the human-generated stories as the reference text, and even for these measures, we see that collaboratively generated text scores do not deviate from the reference score distribution. This analysis indicates that LLMs are able to collaboratively generate stories without any significant changes in quality through sequential prompting. Detailed statistics for each LLM are provided in the Appendix (Table 8).
@@ 4.1 Story Continuity
An important aspect of collaborative storywriting is the notion of continuity. Particularly, we wanted to measure if the different story parts generated sequentially by different LLM authors followed a logical and cohesive plot. To evaluate this, we used a prompt-based evaluation using GPT-4o OpenAI (2024) to discern if two consecutive story parts continue a story well or not. To do this, we develop two types of pairs of story parts: one pair containing the "Correct" or the positive sample which points to the actual next part in the story, or "Incorrect" which is the negative sample that is either a random story part drawn from a different story or from the same one (but not the true next one, any other part). In other words, for the negative samples taken from within the same story, we ensure to not use the correct consecutive part as a negative sample. An example of this setup is illustrated in Table 6, where given a story part, there is an actual continuation ("Correct") and a random negative sample ("Incorrect") story part drawn from either the same story or a different one. The motivation for using the correct and incorrect pairs is to check if the evaluator (GPT-4o) gives high continuity scores to story parts (pairs) that were continuous and low continuity scores to story parts that were not consecutive. This type of A/B testing is a means of demonstrating that our LLM-based evaluation is reasonable/in accordance with expected results. We implicitly account for the biases as follows:
  • Length: we keep the length of story parts approximately similar using a word limit constraint in the generation prompts. This ensures that the lengths of the story parts in every pair being evaluated are similar.
  • Position: The pairs being evaluated are drawn from all the possible positions in the text. In our evaluation, we present a general analysis of how continuous any two consecutive story parts are, irrespective of where they occur in the story, and present an aggregate over all such positions (beginning to end).
  • Self-Bias: We control for self-bias by ensuring that GPT-4o or a similar OpenAI model was not one of the LLMs being used for generation.
Story Part True Continuation Random Negative Sample
Lisa was born to a small village on the outskirts of a vast forest . All her life, she was a regular village girl , happily tending to the family’s crops and tending to any animal that came her way. The locals were fond of her, and she of them. On her eleventh birthday, she received a mysterious gift from the village elder. It was a small, silver orb, engraved with intricate runes. The orb seemed almost sentient, and as she looked upon its surface, a sense of calm washed over her like a gentle breeze. The days that followed were uneventful. Lisa went about her usual duties, always making sure her crops grew strong and healthy. She rescued animals in distress and cared for any sick or injured creatures she found. Life in the village was peaceful and serene. Lisa never questioned why she was different . She trusted her instincts and believed in the good in everyone. The villagers liked her for her pure heart, and her kindness was well-known throughout the land. Yet, as the days turned to years, Lisa couldn’t shake off an inexplicable feeling of restlessness . She spent long hours by the forest , watching the trees sway in the wind, and the sunlight filtering through their leaves. On her eleventh birthday, she received a mysterious gift from the village elder - a beautifully crafted wooden box with intricate carvings of flowers and animals . The box was locked, but the key was inside. Lisa opened it to find an old map of the forest, marked with symbols she had never seen before. Her heart raced as she stared at the map. She knew this was no ordinary gift. The village elder pulled her aside and revealed the truth - Lisa was the chosen one, destined to protect the forest from an ancient evil that had been sleeping for centuries. Lisa was taken aback, but she knew deep down that this was the reason for her restlessness. She spent the next few weeks learning about the forest and the magic that resided within it . The officer stepped forward cautiously, his boots echoing through the otherwise silent corridor . He reached the end, where a single door stood slightly ajar, revealing a flickering light within. Drawing his weapon, he prepared for the unknown dangers that lay beyond. As he pushed the door open wider, he was greeted by a sight that chilled his blood. A room full of masked figures, their eyes glowing with a malevolent light, danced and swayed to an eerie tune. The officer’s heart raced, and he knew that he had stumbled upon a dangerous cult gathering . He tried to retreat, but it was too late. The masked figures noticed him, and with a sinister cackle, they advanced towards him. The officer fired his weapon, hoping to scare them off, but his bullets seemed to have no effect. They closed in on him, their faces twisted into grotesque grins, and the officer knew that he was in grave danger. As the masked figures closed in on him, the officer felt a surge of determination. He would not go down without a fight .
Table 6: To evaluate the continuity of story parts, we sample two potential continuations to a story part, one correct and one incorrect. In this figure, we see an example of a story part on the left and two candidates to be evaluated for continuity. In green are some terms highlighted to show continuity in topics, characters, and plot lines , and in red are highlighted sections of the text that suggest discontinuity .
@@
We then use the following prompt structure to obtain continuity evaluations for both candidate next parts and then compare if the actual continuation was the one with a better continuation evaluation. We provide GPT-4o a pair of texts at a time (out of a total of 600 pairs) and ask it the following question: “Does Part 2 serve as a good continuation of Part 1 in terms of logical flow, coherence, and consistency? Please respond with Yes or No.” GPT-4o generated the binary "yes/no" responses and an explanation. We then compare the percentage of story parts for which either the correct answer got a "Yes" and the incorrect one got a "No" response (Correct Wins), if both candidate continuations were evaluated the same (Neither Wins), or if the Incorrect answer got a "Yes" and the Correct one got a "No". We present the results in Table 5. We see that when the negative story parts are sampled from different stories than the one that the first part belongs to, then there are 0 cases for which the incorrect part wins. This is an expected result, as story parts that belong to different stories would have different topics, plot lines, characters, and so on. Hence, detecting this discontinuity should not be hard, as is the case. An example of a case where Neither Wins and both candidate story parts get an evaluation as being discontinuous is provided in Table 6. As illustrated by this example, it is crucial to note that our prompt encourages the model to evaluate the logical or common-sense flow of the stories. Hence, topic overlap alone was insufficient to mark two parts as continuous. In the harder setting, when the negative sample is also a story part from within the same story, we see that there is a marked increase in both the Neither Wins and Incorrect Wins scenarios (see Table 5) where Incorrect Wins goes from 0 to 9.67% and Neither Wins has an over 30% increase. This is expected since many of the story parts were not the next immediate part but a few sequences apart. They all follow the same story and share some logical flow despite the distance in their occurrences in the story. For example, the fourth part of story might get a positive continuity evaluation, just as the second, third, and fifth would if compared with the first part of the story. Over 85% of Neither Wins story pairs had a continuous evaluation for both candidates (142 out of 166). Thus, out-of-the-box LLMs are able to follow plot lines and logic even when continuing each others’ partially written parts at any point in the story.